What is the smallest value of 9(a^2) + (b^2) +16(c^2) given that 3a-b+4c=16 and

(4/a) - (12/b) +(3/c) = 0


1
Expert's answer
2021-10-08T10:57:36-0400

Given that: 3ab+4c=16(1)3a-b+4c=16\ldots(1)

Also, 4a12b+3c=0\frac{4}{a} - \frac{12}{b} +\frac{3}{c} = 0

4bc12ac+3ababc=04bc12ac+3ab=0(2)\Rightarrow \frac{4bc-12ac+3ab}{abc}=0\\ \Rightarrow 4bc-12ac+3ab=0\ldots(2)

Squaring equation (1),(1), we get:

9a2+b2+16c2+2(3ab4bc+12ac)=09a2+b2+16c22(3ab+4bc12ac)=09a2+b2+16c2=0 [From (2)]9a^2+b^2+16c^2+2(-3ab-4bc+12ac)=0\\ \Rightarrow 9a^2+b^2+16c^2-2(3ab+4bc-12ac)=0\\ \Rightarrow 9a^2+b^2+16c^2=0 \ [From \ (2)]